OpenStudy (anonymous):

Solve: m minus five and nine-tenths equals two and two-fifths. m equals two and one-half m equals three and three-tenths m equals eight and three-tenths m equals nine and one-half

OpenStudy (dangerousjesse):

So you're just asking \(m-5\frac{9}{10}=2\frac{2}{5}\)?

OpenStudy (dangerousjesse):

http://www.wolframalpha.com/input/?i=+m+minus+five+and+nine-tenths+equals+two+and+two-fifths Sorry for sending a link instead of an answer

OpenStudy (anonymous):

what